May 29, 2014

Observer: Zheng Cai, Yuguang Chen

Comments: Theaten clouds went away from 10pm. Clear after 10:30pm.

Time (MST)

10:00 Dome opened.

10:25- 11:10 Smith's high priority target SN2000ch, finish imaging in both g and r, and finish 3x10min spectra, seeing 1.4-1.5''

11:10- 11:25 Telescope re-collimation on stars near Cai's field.

11:25- 11:55 Slew to Cai's field, and acquisition on Cai's mask. It took a while for observers to be confident about the alignment. Main problem is that ModsAlign routine did not correctly calculated the center of two boxes. Observers figured out the reason later. In future, it is more proper to first subtract the bias, and then run ModsAlign on bias-subtracted frame, and this function has been included in the new version of ModsAlign (check with Olga K.)

11:55- 1:50 Finish 5x20min exposure on Cai's mask, seeing varies from 1.2- 1.7, with the average of 1.4''.

1:50- 2:15 The force of the actuators on the mirror exceeded the upper limit. Telescope went to zenith to re-collimate.

2:20- 2:30 acqusition on Smith's high priority target SN2008S. First acquisition failed because of the telescope collimation problem (stars were strongly elongated).

2:30- 2:45 re-collimated the telescope, and re-acquisition on SN2008S.

2:45- 3:55 finish 3x 20min exposure on SN2008S, seeing about 1.3'' on average.

Edited by Zheng Cai

-- %USERSIG{EiichiEgami - 2014-05-26}%

Comments


 
Topic revision: r2 - 30 May 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback